### v1

# Second FRCSyn-onGoing: Winning Solutions and Post-Challenge Analysis to Improve Face Recognition with Synthetic Data

Abstract
Synthetic data is gaining increasing popularity for face recognition technologies, mainly due to the privacy concerns and challenges associated with obtaining real data, including diverse scenarios, quality, and demographic groups, among others. It also offers some advantages over real data, such as the large amount of data that can be generated or the ability to customize it to adapt to specific problem-solving needs. To effectively use such data, face recognition models should also be specifically designed to exploit synthetic data to its fullest potential. In order to promote the proposal of novel Generative AI methods and synthetic data, and investigate the application of synthetic data to better train face recognition systems, we introduce the 2 FRCSyn-onGoing challenge, based on the 2 Face Recognition Challenge in the Era of Synthetic Data (FRCSyn), originally launched at CVPR 2024. This is an ongoing challenge that provides researchers with an accessible platform to benchmark i) the proposal of novel Generative AI methods and synthetic data, and ii) novel face recognition systems that are specifically proposed to take advantage of synthetic data. We focus on exploring the use of synthetic data both individually and in combination with real data to solve current challenges in face recognition such as demographic bias, domain adaptation, and performance constraints in demanding situations, such as age disparities between training and testing, changes in the pose, or occlusions. Very interesting findings are obtained in this second edition, including a direct comparison with the first one, in which synthetic databases were restricted to DCFace and GANDiffFace.
Face biometrics is a very popular area within Computer Vision and Pattern Recognition, finding applications across various domains such as person recognition [1, 2], healthcare [3, 4], and e-learning [5], among others. In recent years, with the fast development of deep learning, significant advances have been made in areas like face recognition (FR) [6, 7], surpassing previous benchmarks. However, FR technology still faces challenges in several research directions, including explainability [8, 9, 10], demographic bias [11, 12], privacy [13, 14, 15], and robustness against adverse conditions [7], such as aging [16], pose variations [17, 18], illumination changes [19], and occlusions [20].
Figure 1: Examples of synthetic identities and variations for different demographic groups using GANDiffFace [21].
Synthetic data has gained popularity as a good solution to mitigate some of these drawbacks [22, 23], allowing the generation of i) a large number of facial images from different non-existent identities, and ii) variability in terms of demographic attributes and scenario conditions. Several Generative AI approaches have been presented in the last couple of years for the synthesis of face images, considering state-of-the-art deep learning methods. Two of the most popular methods are Generative Adversarial Networks (GANs) [24] and Diffusion Models [25]. Popular generative methods can be SynFace [26] and SFace [27], based on GANs, DCFace [28], IDiff-Face [29] and ID3PM [30], based on Diffusion models, GANDiff-Face [21], based on the combination of GANs and Diffusion models, or alternative methods such as DigiFace-1M [31] and ITI-GEN [32], based on computer graphics techniques and a vision-language pre-trained CLIP [33] model, respectively. Some examples of synthetic face images generated using GANDiffFace are shown in Figure 1.
Beyond the generation of synthetic faces, another critical aspect lies in understanding the potential applications and benefits of synthetic data in enhancing FR technology. Recent studies have highlighted a performance gap between FR systems trained only with synthetic data and those trained on real data [28, 26]. Nevertheless, the results achieved in the 1 edition of the Face Recognition Challenge in the Era of Synthetic Data (FRCSyn) [34, 35], emphasize the relevance of synthetic data, either alone or merged with real data, in mitigating challenges in FR, such as demographic bias [34, 35]. Notably, in the 1 FRCSyn-onGoing, only synthetic data from DCFace [28] and GANDiffFace [21] methods were allowed for training FR systems. Additionally, together with novel generative methods, improving FR technology involves refining the design and training processes to address domain gaps between real and synthetic data in certain scenarios. For instance, observations from the 1 FRCSyn-onGoing revealed that most teams considered similar deep learning architectures (e.g., ResNet-100 [36]) and loss functions (e.g., AdaFace [7]), commonly used in FR systems trained with real data.
In order to promote the development of novel face generative methods and the creation of synthetic face databases, as well as investigate the application of synthetic data to better train FR systems, we have organized the 2 FRCSyn-onGoing Challenge, which is based on the 2 FRCSyn Challenge as part of CVPR 2024111https://frcsyn.github.io/CVPR2024.html [37]. In this 2 edition, we introduce new sub-tasks allowing participants to train FR systems using synthetic data generated with their preferred generative frameworks, offering more flexibility compared to the 1 edition [35, 34]. Additionally, new sub-tasks with varied experimental settings are included to explore how FR systems can be trained under both constrained and unconstrained scenarios regarding the amount of synthetic training data. The FRCSyn Challenge aims to address the following research questions:
1.
What are the limitations of FR technology trained only with synthetic data?
2.
Can synthetic data help alleviate current limitations in FR technology?
These questions have become increasingly relevant after the discontinuation of popular real FR databases due to privacy concerns222https://exposing.ai/about/news/ (March, 2024) and the introduction of new regulatory laws333https://artificialintelligenceact.eu (March, 2024).
The foundation of the present article was established in an earlier publication [37], with the current version notably extending it through: i) a more extensive description and analysis of the top synthetic face generation methods and FR systems presented so far in this 2 FRCSyn-onGoing, including key graphical representations of the proposed systems to improve the understanding of the reader, ii) incorporating additional metrics in the evaluation of the proposed FR systems in order to analyze different operational scenarios, iii) presenting an in-depth analysis of the performance achieved for various demographic groups and databases used for evaluation, together with novel figures and tables, and iv) a direct comparison between the results obtained in this 2 edition and the ones obtained in 1 edition [34, 35], highlighting very interesting findings.
The remainder of the article is organized as follows. Section II describes the databases considered at the 2 FRCSyn-onGoing. Section III explains the experimental setup of the challenge, including the different tasks and sub-tasks, the experimental protocol, metrics, and restrictions. In Section IV, we describe the approaches proposed by the top-6 participating teams. Section V presents the best results achieved so far in the different tasks and sub-tasks of 2 FRCSyn-onGoing, emphasizing the key results of the challenge. Finally, in Section VI, we provide some conclusions, highlighting potential future research directions in the field.
II Second FRCSyn-onGoing: Databases
II-ASynthetic Databases
One of the main novelties of the 2 FRCSyn-onGoing is that there are no restrictions in terms of the generative methods used to create synthetic data. Unlike the 1 FRCSyn-onGoing, where only synthetic data created using DCFace [28] and GANDiffFace [21] was available, in this 2 edition we allow participants to use any generative framework of their choice to create synthetic data, limiting in some sub-tasks the number of synthetic face images used to train the FR systems (more details in Section III-A). As a reference, after the registration in the challenge, we provide all the participants with a list of possible state-of-the-art generative frameworks. For completeness, we summarize next and in Table I the most popular approaches available at the beginning of the challenge:
TABLE I: Description of some possible generative methods and synthetic databases that can be used by participants in the 2 FRCSyn-onGoing.
Id = Identities, Img = Images
DCFace444https://github.com/mk-minchul/dcface [28]: This framework is entirely based on Diffusion models, composed of a sampling stage for the generation of synthetic identities , and a mixing stage for the generation of images with the same identities from the sampling stage and the style selected from a “style bank” of images .
GANDiffFace555https://github.com/PietroMelzi/GANDiffFace [21]: This framework combines StyleGAN [38] and a Diffusion Model, i.e., DreamBooth [39], to generate fully synthetic FR databases with desired properties such as human face realism, controllable demographic distributions, and realistic intra-class variations (e.g., changes in pose, expression, and occlusions). Graphical examples are shown in Figure 1.
IDiff-Face666https://github.com/fdbtrs/IDiff-Face [29]: This framework uses a Diffusion Model conditioned on identity context, which allows the model to either generate variations of existing authentic images by using authentic embeddings or to generate novel synthetic identities by using synthetic face embeddings. The authors presented two distinct datasets: one by generating identity context in a two-stage process, and the other through a synthetic uniform representation.
DigiFace-1M777https://github.com/microsoft/DigiFace1M [31]: This framework can generate large-scale synthetic face images with many unique subjects based on 3D parametric model rendering. It considers the method introduced by Wood et al. [40], tackling the ethical and labeling problems associated with the generation of synthetic data.
ID3PM [30]: This framework considers a Diffusion Model to perform an inversion of a FR model generating new images from Gaussian noise with various backgrounds, lighting, poses, and expressions while preserving the identity.
SYNFace999https://github.com/haibo-qiu/SynFace [26]: This framework uses DiscoFaceGAN [41] to generate face images with different identities from a Mixup Face Generator.
ITI-GEN101010https://github.com/humansensinglab/ITI-GEN[32]: This framework uses CLIP [33] to generate embeddings to translate the visual attribute differences into natural language differences and perform a Text-to-Image generation that is inclusive.
These are just some possible generative frameworks, with the corresponding synthetic databases available, that can be used by participants. But, as indicated before, the purpose of the 2 FRCSyn-onGoing is to promote the proposal of novel generative methods and the creation of better synthetic databases to improve the performance of FR systems. It is important to mention that in the 2 FRCSyn-onGoing, synthetic data is exclusively used in the training stage of FR technology, replicating realistic operational scenarios.
II-BReal Databases
TABLE II: Description of the real databases considered for evaluation in the 2 FRCSyn-onGoing.
Id = Identities, Img = Images
For the training of the FR systems participants are allowed to use only the CASIA-WebFace [42] as real data (depending on the sub-task, please see Section III-A). This database contains face images of real identities collected from the web. For the final evaluation of the proposed FR systems, we consider the same four real databases used at the 1 FRCSyn Challenge [34, 35], as they consider key challenges in FR such as demographic bias, pose variations, aging, and occlusions. We summarize next and in Table II each of them:
BUPT-BalancedFace [43] is designed to address performance disparities across different ethnic groups. We relabel it according to the FairFace classifier [47], which provides labels for ethnicity (White, Black, Asian, Indian) and gender (Male, Female). We then consider the eight demographic groups obtained from all possible combinations of four ethnic groups and genders. We are aware that these groups do not comprehensively represent the entire spectrum of real world ethnic diversity. Nevertheless, the selection of these categories, while imperfect, is primarily driven by the need to align with the demographic categorizations used in BUPT-BalancedFace to facilitate easier and more consistent evaluation.
AgeDB [44] contains facial images featuring the same subjects at different ages in different environmental contexts.
CFP-FP [45] presents facial images from subjects with great changes in pose, with both frontal and profile images, and different environmental contexts.
ROF [46] consists of occluded faces with both upper face occlusion, due to sunglasses, and lower face occlusion, due to masks.
Finally, it is important to highlight that, as different databases are considered for training and evaluation, we also intend to analyze the generalization ability of the proposed FR systems.
TABLE III: Tasks and sub-tasks for the 2 FRCSyn-onGoing and their respective metrics and databases.
SD = Standard Deviation. ID = Identity
Task 1: synthetic data for demographic bias mitigation
Sub-Task 1.2: [unconstrained] training exclusively with synthetic data
Train: no restrictions in terms of the number of face images.
Eval: BUPT-BalancedFace.
Sub-Task 1.3: [constrained] training with real and synthetic data
Train: CASIA-WebFace, and maximum face synthetic images.
Eval: BUPT-BalancedFace.
Task 2: synthetic data for overall performance improvement
Baseline: training with only CASIA-WebFace.
Metrics: accuracy (for each evaluation database).
Ranking: average accuracy, see Section III-C for more details.
Sub-Task 2.1: [constrained] training with only synthetic data
Train: maximum face images.
Eval: BUPT-BalancedFace, AgeDB [44], CFP-FP [45], and ROF [46].
Sub-Task 2.2: [unconstrained] training with only synthetic data
Train: no restrictions in terms of the number of face images.
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF.
Sub-Task 2.3: [constrained] training with real and synthetic data
Train: CASIA-WebFace, and maximum face synthetic images.
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF.
III Second FRCSyn-onGoing: Setup
Due to the success of the 1 FRCSyn-onGoing [34, 35], we also decided to run the 2 edition in Codalab111111https://codalab.lisn.upsaclay.fr/competitions/16970, an open-source framework designed for conducting scientific competitions and benchmarks. On this platform, participants can find the competition’s requirements and limitations and can submit their scores to automatically obtain i) the evaluation metrics of their system, and ii) the position on the challenge leaderboard.
III-ATasks
Similar to the 1 FRCSyn-onGoing [35, 34], in this 2 edition we also explore the application of synthetic data for training FR systems, with a specific focus on addressing two critical aspects in current FR technology: i) mitigating demographic bias, and ii) enhancing overall performance under challenging conditions that include variations in age and pose, the presence of occlusions, and diverse demographic groups. To investigate these two areas, we consider two different tasks, each comprising three sub-tasks. Each sub-task considers different types (real/synthetic) and amounts of data for training the FR systems. Consequently, the 2 edition comprises 6 different sub-tasks. We summarize in Table III the key aspects of the experimental protocol, metrics, and restrictions for each sub-task.
Task 1
The first task focuses on using synthetic data to mitigate demographic biases within FR systems. To evaluate the performance of these systems, we create sets of mated and non-mated comparisons using subjects from the BUPT-BalancedFace database [43]. We consider the eight demographic groups defined in Section II-B, which result from the combination of four ethnicities (White, Black, Asian, and Indian) and two genders (Male and Female), ensuring a balanced representation across these groups in the comparison lists. For non-mated comparisons, we exclusively pair subjects within the same demographic group, as these hold greater relevance compared to non-mated comparisons involving subjects from different demographic groups.
Task 2
The second proposed task focuses on using synthetic data to enhance the overall performance of FR systems under challenging conditions. To assess the effectiveness of the proposed systems, we use lists of mated and non-mated comparisons selected from subjects from the different evaluation databases, each one designed to address specific challenges in FR. Specifically, BUPT-BalancedFace is used to consider diverse demographic groups, whereas AgeDB, CFP-FP, and ROF to assess age, pose, and occlusion challenges respectively.
III-BExperimental protocol
Training
The six sub-tasks introduced in the 2 FRCSyn-onGoing are mutually independent. This implies that participants have the flexibility to participate in any number of sub-tasks based on their preferences. For each selected sub-task, participants are required to develop a FR system and train it twice: i) using the authorized real databases exclusively, i.e., CASIA-WebFace [42], and ii) following the specific requirements of the chosen sub-task, as summarized in Table III. According to this protocol, participants must provide both the baseline system and the proposed system for each specific sub-task. The baseline system plays a critical role in evaluating the impact of synthetic data on training and serves as a reference point for comparing the proposed model against the conventional practice of training only with real databases. To maintain consistency, the baseline FR system, trained exclusively with real data, and the proposed FR system, trained according to the specifications of the selected sub-task, must have the same architecture and training protocol.
Evaluation
In each sub-task, participants received the comparison files comprising both mated and non-mated comparisons, which are used to evaluate the performance of their proposed FR systems. Task 1 involves a single comparison file containing balanced comparisons of different demographic groups of the BUPT [43] database, while Task 2 comprises four comparison files, each corresponding to every specific real-world databases considered (i.e., BUPT, AgeDB [44], CFP-FP [45], and ROF [46]). During the evaluation of each sub-task, participants are required to submit two files per database through the Codalab platform: i) the scores of the baseline system, and ii) the scores of the proposed system. Finally, for each sub-task, participants must submit a file including the decision threshold for each FR system (i.e., baseline and proposed). The submitted scores must fall within the range of , with lower scores indicating non-mated comparisons, and vice versa.
III-CEvaluation Metrics
We evaluate the FR systems using a protocol based on lists of mated and non-mated comparisons for each sub-task and database. From the scores and thresholds provided by participants, we calculate the binary decision and the verification accuracy. Additionally, we calculate the gap to real (GAP) [28] as follows: , with REAL representing the verification accuracy of the baseline system and SYN the verification accuracy of the proposed system, trained with synthetic (or real + synthetic) data. Other metrics such as False Non-Match Rate (FNMR) at a fixed operational point, or the Area Under the ROC Curve which are very popular for the analysis of FR systems in real-world applications, are also computed from the scores provided by participants. Next, we explain how participants are ranked in the different tasks.
Task 1
To rank participants and determine the winners of Sub-Tasks 1.1, 1.2, and 1.3, we closely examine the trade-off between the average (AVG) and standard deviation (SD) of the verification accuracy across the eight demographic groups defined in Section II-B. We define the trade-off metric (TO) as follows: . This metric involves plotting the average accuracy on the x-axis and the standard deviation on the y-axis in a 2D space. Multiple 45-degree parallel lines are drawn to identify the winning team, whose performance is located on the far right of these lines. With this proposed metric, we reward FR systems that achieve good levels of performance and fairness simultaneously, unlike common benchmarks based only on recognition performance. The standard deviation of verification accuracy across demographic groups is a common metric for assessing bias and should be reported by any work addressing demographic bias mitigation.
Task 2
To rank participants and establish the winners in Sub-Tasks 2.1, 2.2, and 2.3, we examine the average verification accuracy from the four different databases designated for evaluation, as described in Section II. This approach enables us to assess four main challenges of FR technologies: i) the representation of diverse demographic groups, ii) the impact of aging on recognition, iii) the variations in facial pose, and iv) the challenges made by occlusions. This evaluation provides a comprehensive overview of FR systems in real operational scenarios.
III-DRestrictions
Participants have the freedom to choose the FR system for each task as long as the number of Floating Point Operations Per Second (FLOPs) of the system does not exceed 50 GFLOPs. This threshold has been established to facilitate the exploration of innovative architectures and encourage the use of diverse models while preventing the dominance of excessively large models. Participants are also free to use their preferred training modality, with the requirement that only the specified databases are used for training. Generative models cannot be used to generate supplementary data. Participants are allowed to use non-face databases for pre-training purposes and use traditional data augmentation techniques using the authorized training databases. To maintain the integrity of the evaluation process, the organizers reserve the right to disqualify participants if anomalous results are detected or if participants fail to adhere to the challenge’s rules.
IV Second FRCSyn-onGoing: Systems Description
In this 2 FRCSyn-onGoing, we encourage participants to propose novel Generative AI methods for the creation of synthetic data. Besides, we also give participants the freedom to choose the FR architecture and training methods. Table LABEL:tab:data_meth summarizes for each team the key information in terms of the proposed synthetic data and FR system. Teams are arranged by their average ranking in the 6 sub-tasks from the 2 FRCSyn-onGoing. In general, we can see that most teams have decided to use synthetic data from DCFace [28] and IDiff-Face [29] databases, improving also the original data through cleaning and selection approaches, among other more sophisticated techniques. Also, regarding the FR technologies, most of them are based on ResNet [36] and IResNet [48] architectures, with AdaFace [7] and ArcFace [6] as the main used losses. However, some teams proposed their own methods to generate synthetic facial images, as well as to train their FR models. Next, we describe the specific details of the top-11 proposed systems in the 2 FRCSyn-onGoing.
ADMIS (All sub-tasks)
This team comprises members from Fudan University and Tencent Youtu Lab, China. They used a Latent Diffusion Model (LDM) [49] based on IDiff-Face [29] to synthesize faces. The LDM was conditioned using identity embeddings as contexts, extracted from faces by a pretrained ElasticFace [50] model with an IResNet-101 [48] backbone. They trained the LDM with the CASIA-WebFace [42] database. As the LDM takes the ID embeddings as context, they considered an unconditional Denoising Diffusion Probabilistic Model (DDPM) trained on the FFHQ database [51] as a context generator. Specifically, they used the DDPM to generate faces with arbitrary identities, known as context faces. They exploited the same ElasticFace model to extract the embeddings from the context faces. To encourage the quality and the distinctiveness of identity from later LDM-generated faces, they filtered the embeddings by setting a minimum cosine similarity threshold of between arbitrary pairs of embeddings. This yields embeddings with discriminative identities. Furthermore, they accelerated the sampling process of the LDM by Denoising Diffusion Implicit Models (DDIM) [52]. For the training of the FR model, they generated images for each context. They adopted the ID oversampling strategy from DCFace [28] and performed it five times for each ID to enhance consistency. As a result, contexts were used for Sub-Tasks 1.1 and 2.1, while for Sub-Tasks 1.2 and 2.2. For Sub-Tasks 1.3 and 2.3, they expanded Sub-Tasks 1.1 and 2.1 with the CASIA-WebFace database. Both the baseline and proposed FR models used IResNet-101 architectures. They applied the ArcFace [6] loss with a batch size of and an initial learning rate of for epochs. The learning rate was divided by at epochs , , and . They also used random cropping augmentation during training. Their proposed architecture is described in Figure 2.
Figure 2: Framework proposed by the ADMIS team.Figure 3: Framework proposed by the OPDAI team.Figure 4: Framework proposed by the ID R&D team.
OPDAI (All sub-tasks)
This team comprises members from Interactive Entertainment Group of Netease Inc, China. They initially used the DCFace [28] database, generating then more face images for each ID with large pose variations and occlusions using Photomaker [53]. They randomly replaced these images in the original DCFace data to ensure that the total number of samples meets the requirement of . During the Photomaker inference, they adopted a batch size of and used random prompts including age, pose, and image quality to ensure the diversity of the generated samples. For Sub-Tasks 1.2 and 2.2, they combined this data with the version of DCFace, while for Sub-Tasks 1.3 and 2.3, it was merged with CASIA-WebFace [42]. For Sub-Tasks 1.2, 1.3, 2.2, and 2.3 they did not merge nor denoise samples from different databases, following the Partial FC approach [64]. Regarding the FR model, they obtained the loss of different databases in independent AdaFace [7] heads, calculating the final loss as the average of the multiple heads. Both baseline and proposed models are based on IResNet-100 [48] architectures, with horizontal flipping. Their proposed architecture is described in Figure 3.
This team comprises members from ID R&D Inc, USA. To generate the synthetic data, they used two models trained on WebFace42M [65]. The first model was based on Hourglass Diffusion Transformers (HDT) [54]. It was trained focusing on conditional flow matching [66] and following the classifier-free guidance approach [67]. Identity embeddings were used directly, whereas style embeddings were processed through a Vector Quantised-Variational AutoEncoder (VQVAE) [55]. Specifically, for head position, embeddings were allocated for VQVAE processing, while age and facial expression attributes were represented with embeddings each. During inference, combinations of these embeddings were randomly selected. The second model used to generate synthetic data was based on StyleNAT [56], enhanced with a FR model. This network was trained using auxiliary sources of supervision: a pre-trained FR network with Prototype Memory [68] (for identity supervision) and a pre-trained face attribute classification network (for style supervision). To create their synthetic data, they used classifier weights of the trained Prototype Memory to get identity embeddings, of which were randomly selected and were uniformly sampled from the clusters obtained using k-means, to get demographic diversity. For each identity, they generated face images using each of the two generative models. In the first stage, identity embeddings were used by the HDT model to get images for each identity. Then of these images were included in the training dataset. Identity and style embeddings were taken from the remaining images and used as a condition to generate different images with the StyleNAT model. These images were also put into the training dataset. Regarding the FR model, they trained an IResNet-200 [48] with UniFace [58] loss for epochs. One network was trained with color, geometric augmentations, and FaceMix-B [57], and the other one using only random horizontal flipping. These two networks were combined in an “ensemble”, where the first one received the original image, and the second one a mirrored copy. They used the same model for Sub-Tasks 1.1, 1.2, 2.1 and 2.2. For Sub-Tasks 1.3 and 2.3, they combined the synthetic data with the CASIA-WebFace [42] data, training two models, one on the mixed data, and the other on the CASIA-WebFace. Their proposed architecture is described in Figure 4.
Figure 5: Framework proposed by the K-IBS-DS team.
K-IBS-DS (All sub-tasks)
This team comprises members from Korea Advanced Institute of Science & Technology and Institute for Basic Science, South Korea. They used DCFace [28] with and face images (depending on the sub-task). Regarding the FR model, they used several IResNet [48] models of , , and layers with Squeeze-and-Excitation (SE) blocks [60]. Inspired by SlackedFace [59], they made two modifications to enhance the AdaFace [7] FR classifier: i) they used renormalized uniform initialization as a more reliable weight initialization for uniformity across identity prototypes in the unit sphere and ii) replaced the L2-norm with the powered-norm (p-norm), or face recognizability index from [59], which integrates the L2-norm with the learned embedding proximity. The training stage was in line with [28] and [31], including optimizer, learning rate, etc. For Sub-Tasks 1.3 and 2.3, the first subjects of the CASIA-WebFace [42] were assigned for training, and the remaining ones for performance validation using random pairs with challenging conditions (identified based on the poorest L2-norm values [21]). The final score was obtained by aggregating the comparison scores of the different IResNet models, along with the horizontally flipped instances through score fusion. All training and test sets were realigned using RetinaFace [69], followed by a similarity transformation. For all sub-tasks, aggressive data augmentations were applied, including random horizontal flipping, photometric operations, cropping, resizing, and the addition of sunglasses and masks. Their proposed architecture is described in Figure 5.
Figure 6: Framework proposed by the Idiap-SynthDistill team.
CTAI (All sub-tasks)
This team comprises members from China Telecom AI, China. By analyzing popular synthetic data, they found that intra-class and inter-class noise was widely present. Data cleaning can effectively remove the bad examples of synthetic data and retain important images from a large amount of synthetic data. In order to select the optimal synthetic data, they first trained an IResNet-100 [48] model with SE blocks [60] using CASIA-WebFace [42] to extract features of synthetic images from DCFace [28], GANDiffFace [21], and DigiFace-1M [31]. Subsequently, they used DBSCAN clustering to segregate intra-class noise and removed IDs with a class center feature cosine similarity greater than . Finally, they used the cleaned synthetic data merged with CASIA-WebFace to finetune the IResNet-100 for a second data refinement. From the final refined synthetic dataset, they sampled face images while retaining as many IDs as possible to build their synthetic training set. Regarding the FR model, in particular Sub-Task 2.3 in which they achieved their highest position among all sub-tasks, they trained IResNet-100 with AdaFace loss [7] (A1) and CosFace loss [61] (A2) with mask and occlusion augmentation on CASIA-WebFace and the refined synthetic data. They used an ensemble of A1, A2, and a model trained with only synthetic data. Furthermore, data augmentation was considered to enhance all features.
This team comprises members from Idiap Research Institute, EPFL, and Université de Lausanne, Switzerland. The proposed method was based on SynthDistill [62], which is an end-to-end approach, generating synthetic images and training the FR model in the same training loop. Instead of using the pre-trained model in a separate step, they directly used it in the training loop for supervision, while a new student FR model was trained fully using synthetic data generated from a StyleGAN model [51]. For generating synthetic images, they trained StyleGAN2 [70] with the CASIA-WebFace database [42] and then dynamically generated 20 synthetic images during training based on the training loss. For the dynamic image generation, they used the training loss from every iteration as feedback to find the most difficult synthetic image in each batch and then re-sampled a new batch of synthetic images in the intermediate latent space of StyleGAN near the latent vector of the most difficult sample. If the loss value was high, they re-sampled with a relatively small standard deviation around the difficult sample and generated similar images, but if the loss value was small they re-sampled with a higher standard deviation, generating images with more variations. Throughout the process, the generated images were resized to before being fed to the FR models. They used a pre-trained FR model with IResNet-101 [48] architecture trained with CosFace [61] loss on a subset of the WebFace260M database [65] and trained a new model as a student network with the same architecture using synthetic data with their dynamic synthetic image generation approach. They used the Adam optimizer with an initial learning rate of and trained their student model with the same loss function as in [62]. For thresholding, a subset of DCFace [28] was used to determine the optimal threshold for maximizing verification accuracy, using a 10-fold cross-validation approach based on a random selection of identities and comparison pairs. Their proposed architecture is described in Figure 6.
Figure 7: Architecture proposed by the INESC-IGD team.
INESC-IGD (All sub-tasks)
This team comprises members from INESC TEC and Universidade do Porto, Portugal, and Fraunhofer IGD, Germany. For the training dataset, they merged DCFace [28], IDiff-Face Uniform, and IDiff-Face Two-stage [29] databases and then labeled the data with ethnicity labels using a similar approach to [71]. For Sub-Tasks 1.1 and 2.1, they created a synthetic training dataset containing face images by sampling balanced identities, in terms of ethnicity labels. For Sub-Tasks 1.2 and 2.2, they created a synthetic training dataset containing face images by sampling identities from the training datasets. For Sub-Tasks 1.3 and 2.3, two instances of ResNet-100 [36] were trained, one on CASIA-WebFace [42] and the other on a subset of synthetic datasets (e.g., images of identities). For all sub-tasks, they trained a ResNet-100 with ElasticCosFac-Plus loss [50] using the settings presented in [50]. During the testing phase of Sub-Tasks 1.3 and 2.3, feature embeddings were obtained from trained models and the weighted sum of score-level fusion was used. During the FR training of all sub-tasks, the training datasets were augmented using RandAug and occluded augmentation [72] with probabilities of . The occluded augmentation followed protocol 4 proposed in [72], leading to occlusions on the eyes, lower face, upper face, or a combination of the eyes occlusion with the others. Occluded augmentations boosted the performance, as synthetic data has a lower frequency of natural occlusions such as beard and makeup [73]. Their proposed architecture is described in Figure 7.
This team comprises members from Fraunhofer IGD, Germany, University of Cagliari, Italy, and University of Ljubljana, Slovenia. They used the DCFace [28] synthetic database as it led to remarkable performance gains under well-known evaluation benchmarks for face verification, while combined with real data [74]. Also, they considered synthetic data generated with IDiff-Face [29] and ExFaceGAN [63] in Sub-Tasks 1.1, 1.2, 2.1, and 2.2. The ExFaceGAN data was generated using an identity disentanglement approach on pretrained GAN-Control [75]. Regarding the FR model, they trained a ResNet-100 [36] network using CosFace loss [61] with a margin penalty of 0.35 and a scale term of 64. The similarity mean difference between real-only and synthetic-only samples was scaled and added to the loss value. They trained the FR model for epochs with a batch size of and an initial learning rate of , which was divided by after , , , and epochs. During the training phase, the synthetic samples were augmented using RandAugment with operations and a magnitude of 16, following [76, 74]. For Sub-Tasks 1.3 and 2.3, the selected synthetic dataset was combined with CASIA-Webface [42], obtaining a total of images from identities. Their proposed architecture is described in Figure 8.
Figure 8: Framework proposed by the UNICA-IGD-LSI team.
SRCN_AIVL (Sub-Task 1.1)
This team comprises members from Samsung Electronics (China) R&D Centre, University of Science and Technology, IIE, CAS, and MAIS, CASIA, China. They selected samples from the DCFace [28] database and labeled the ethnicity of each subject, as the racial distribution gap may lead to bad performance in testing. Based on this approach, they trained IDiff-Face [29] with CASIA-WebFace [42] database generating synthetic face images of specific races. Regarding the FR system, they used two custom ResNet-101 [36] trained with AdaFace loss [7] function. The models were trained for epochs with an initial learning rate of and a batch size of , which was adjusted at predefined milestones. Their training data underwent further preprocessing, including padding crop augmentation, low-resolution augmentation, photometric augmentation, random grayscale, and normalization. The threshold was determined by the 10-fold optimal threshold in the validation set. For the inference, data preprocessing involved an MTCNN [77] for the alignment and resizing all data to . After cropping and alignment, they fed the image and the flipped image into the two models. Finally, after obtaining the two feature embeddings, they combined them and performed the similarity calculation with these embeddings. Their proposed architecture is described in Figure 9.
Figure 9: Framework proposed by the SRCN_AIVL team.
CBSR-Samsung (Sub-Tasks 1.3 and 2.3)
This team comprises members from Samsung Electronics (China) R&D Centre, IIE, CAS, and MAIS, CASIA, China. They first trained a FR model using CASIA-WebFace [42]. Then, they used it to de-overlap DCFace [28] from CASIA, as DCFace was trained using that real database. For the synthetic dataset, they compared the performance of models trained with three synthetic databases, including GANDiffFace [21], DCFace, and IDiff-Face [29], and finally selected DCFace as the only synthetic training set. They created a validation dataset including three subsets for three different testing scenarios: i) random sample pairs from DCFace, simulating age variability and demographic groups as in AgeDB [44] and BUPT-Balanced [43] databases, respectively; ii) randomly positioned vertical bar masks to the images to simulate the self-occlusion due to as considered in CFP-FP database [45]; and iii) add mask and sunglasses to the images by detecting the landmarks via FaceX-Zoo [78], simulating the ROF database [46]. This is done following [79]. All validation subsets consist of positive pairs and negative pairs. Finally, they concatenated these subsets as the validation set. Subsequently, they conducted an intra-class clustering for all datasets using DBSCAN ( threshold) and removed the samples that were separated from the class center. Regarding the FR model, they merged the refined datasets and trained IResNet-100 [48] with AdaFace loss [7]. In addition, they adopted two augmentation strategies, i.e., photometric augmentation and rescaling. After that, they trained two FR models using occlusion augmentation with and probability, respectively. Finally, they submitted the average similarity score of the two models. Their proposed architecture is shown in Figure 10.
BOVIFOCR-UFPR (Sub-Tasks 1.2 and 2.1)
This team comprises members from Federal University of Paraná, Federal Institute of Mato Grosso, and unico - idTech, Brazil. They chose DCFace [28] as the synthetic database and randomly removed identities with images per ID to reduce the number to follow the rules. For the FR model, they used a ResNet-100 [36] as the backbone, trained with the ArcFace [6] loss function. The images used for training were augmented using a Random Flip with a probability of . They also applied random erasing and RandAugment as additional augmentations. To validate their model they subsampled images from DCFace, generating genuine and impostor pairs, and used these pairs to select the best threshold to classify the proposed model output scores. Their proposed architecture is described in Figure 11.
Figure 10: Framework proposed by the CBSR-Samsung team.Figure 11: Framework proposed by the BOVIFOCR-UFPR team.
TABLE IV: Results of the teams that ranked among the top-6 in at least one sub-task, ordered by the average rank in all the sub-tasks. For each team, we report the ranking metric and the position across all the sub-tasks. The best result of each sub-task is highlighted in bold. We mark with a ‘-’ if the team did not participate in a sub-task. TO = Trade-off, AVG = average accuracy
Task 1: Bias Mitigation
Task 2: Overall Improvement
Team
Task 1.1
TO [%]
Task 1.2
TO [%]
Task 1.3
TO [%]
Task 2.1
AVG [%]
Task 2.2
AVG [%]
Task 2.3
AVG [%]
ADMIS
94.30 (2)
95.72 (2)
96.50 (1)
91.19 (3)
92.92 (2)
94.15 (5)
OPDAI
93.75 (4)
94.12 (3)
95.96 (4)
91.93 (1)
92.04 (3)
95.23 (2)
ID R&D
96.73 (1)
96.73 (1)
86.73 (8)
91.86 (2)
91.86 (4)
94.05 (6)
K-IBS-DS
92.91 (6)
93.72 (5)
96.17 (2)
91.05 (4)
91.61 (5)
95.42 (1)
CTAI
93.21 (5)
93.21 (6)
95.41 (7)
90.59 (5)
90.59 (6)
94.56 (3)
Idiap-SynthDistill
89.70 (9)
93.50 (1)
INESC-IGD
92.28 (7)
94.05 (4)
95.65 (5)
83.16 (8)
85.40 (7)
89.43 (8)
UNICA-IGD-LSI
91.89 (8)
91.89 (7)
96.00 (3)
87.80 (7)
87.80 (8)
92.79 (7)
SRCN_AIVL
94.06 (3)
CBSR-Samsung
95.57 (6)
94.20 (4)
BOVIFOCR-UFPR
90.48 (8)
89.97 (6)
V Second FRCSyn-onGoing: Results
Next, we describe in Sections V-A and V-B the main results achieved in Tasks 1 and 2, respectively. These results are further analyzed in Section V-C focusing on specific demographic groups and individual databases. Finally, we discuss in Section V-D common trends among the different teams and compare the results with those obtained in the 1 edition. We present in Table IV the ranking and key results of the 2 FRCSyn-onGoing.
V-ATask 1: Bias Mitigation
Table V shows the results achieved by participants in Task 1, focused on demographic bias mitigation. Teams are ranked by descending order of TO, which tends to correlate to the ascending order of SD (i.e., from less to more biased FR systems). Notably, the winner of Sub-Tasks 1.1 and 1.2, ID R&D (96.73% TO), demonstrates a significant negative GAP value (-5.31%), showing a higher performance when training the FR system with synthetic data compared to real data (i.e., CASIA-WebFace [42]). Furthermore, in Sub-Task 1.1, we can observe that teams with negative GAP values considered Diffusion Models for the generation of synthetic data (i.e., ID R&D uses HDT, SRCN_AIVL combines DCFace and IDiff-Face, and CTAI combines DCFace and GANDiffFace), showing that this generation method may work better than real data in scenarios with limited data. Next, after removing the limitation on the number of synthetic images (i.e., Sub-Task 1.2), the TO value of most FR systems increases, which leads to performance and fairness improvement simultaneously. For instance, for the ADMIS team (ranked top-2 in both Sub-Task 1.1 and 1.2), the TO value increases to 95.72% in Sub-Task 1.2 (i.e., 1.42% TO improvement compared to Sub-Task 1.1). Also, the GAP value decreases from 1.47% to -0.56%, obtaining better results when increasing the amount of synthetic data in comparison to limited real data (i.e., CASIA-WebFace). Another example is the OPDAI team, which raised from top-4 to top-3 positions between Sub-Task 1.1 and 1.2. Its TO value increases to 94.12% (i.e., 0.37% TO improvement from Sub-Tasks 1.1 to 1.2), and the GAP value is reduced from 1.02% to 0.71%. These findings emphasize the potential of generating large number of synthetic face images from different demographic groups to mitigate bias in existing FR technology. Finally, we analyze in Sub-Task 1.3 the case of using both, real and synthetic data, in the FR training process. In general, we can observe considerable improvements in terms of TO values, along with higher negative GAP values for all the top-6 teams, e.g., ADMIS (96.50% TO, -1.33 GAP), K-IBS-DS (96.17% TO, -1.37% GAP), and UNICA-IGD-LSI (96.00% TO, -5.33% GAP). These results highlight that the combination of synthetic and real data achieves superior FR performance compared to the case of only training with real data. Finally, it is noteworthy to compare the best results achieved in Sub-Task 1.2, i.e., unconstrained synthetic data, and Sub-Task 1.3, i.e., constrained synthetic + real data. The ID R&D team achieves 96.73% TO in Sub-Task 1.2, whereas ADMIS achieves 96.50% TO in Sub-Task 1.3, showing that unlimited synthetic data for training can even outperform FR systems trained with limited synthetic + real data. These results motivate the use of synthetic data for demographic bias mitigation, improving at the same time privacy as no real identities are seen by the network.
TABLE V: Ranking for the three sub-tasks considered in Task 1. For each sub-task, we highlight in bold the best team according to the Trade-Off. TO = Trade-Off, AVG = Average accuracy, SD = Standard Deviation of accuracy, FNMR = False Non-Match Rate, FMR = False Match Rate, AUC = Area Under Curve, GAP = Gap to Real.
Sub-Task 1.1 (Bias Mitigation): Synthetic Data (Constrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ID R&D
96.73
97.55
0.82
3.17 (1)
99.57 (1)
-5.31
2
ADMIS
94.30
95.10
0.80
11.38 (3)
98.96 (3)
1.47
3
SRCN_AIVL
94.06
95.12
1.07
10.72 (2)
98.83 (4)
-0.54
4
OPDAI
93.75
94.92
1.17
11.85 (4)
99.51 (2)
1.02
5
CTAI
93.21
94.74
1.53
14.38 (5)
98.33 (6)
-0.63
6
K-IBS-DS
92.91
94.11
1.2
15.03 (6)
98.47 (5)
1.58
Sub-Task 1.2 (Bias Mitigation): Synthetic Data (Unconstrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ID R&D
96.73
97.55
0.82
3.17 (1)
99.57 (2)
-5.31
2
ADMIS
95.72
96.50
0.78
6.33 (2)
99.51 (3)
-0.56
3
OPDAI
94.12
95.22
1.11
10.78 (3)
98.92 (4)
0.71
4
INESC-IGD
94.05
95.22
1.17
11.03 (4)
98.70 (5)
1.04
5
K-IBS-DS
93.72
94.88
1.16
12.75 (5)
99.66 (1)
0.77
6
CTAI
93.21
94.74
1.53
14.38 (6)
98.33 (6)
-0.63
Sub-Task 1.3 (Bias Mitigation): Synthetic + Real Data (Constrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ADMIS
96.50
97.25
0.75
3.90 (1)
99.72 (1)
-1.33
2
K-IBS-DS
96.17
96.92
0.75
5.88 (4)
99.54 (2)
-1.37
3
UNICA-IGD-LSI
96.00
96.70
0.70
5.90 (5)
99.49 (3)
-5.33
4
OPDAI
95.96
96.80
0.84
4.90 (2)
99.54 (2)
-0.03
5
INESC-IGD
95.65
96.33
0.67
6.15 (6)
99.18 (5)
-0.12
6
CBSR-Samsung
95.57
96.54
0.97
5.00 (3)
99.41 (4)
-24.43
TABLE VI: Ranking for the three sub-tasks considered in Task 2. For each sub-task, we highlight in bold the best team according to the Average Accuracy. AVG = Average accuracy, FNMR = False Non-Match Rate, FMR = False Match Rate, AUC = Area Under Curve, GAP = Gap to Real.
Sub-Task 2.1 (Overall Improvement): Synthetic Data (Constrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
OPDAI
91.93
17.63 (2)
97.30 (2)
3.09
2
ID R&D
91.86
10.36 (1)
97.48 (1)
2.99
3
ADMIS
91.19
20.41 (3)
97.04 (3)
2.78
4
K-IBS-DS
91.05
24.87 (6)
96.09 (6)
2.60
5
CTAI
90.59
21.88 (4)
96.40 (5)
-1.94
6
BOVIFOCR-UFPR
89.97
24.04 (5)
96.70 (4)
3.71
Sub-Task 2.2 (Overall Improvement): Synthetic Data (Unconstrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
Idiap-SynthDistill
93.50
9.17 (1)
97.17 (4)
-0.05
2
ADMIS
92.92
14.45 (3)
97.76 (1)
0.21
3
OPDAI
92.04
16.48 (4)
97.41 (3)
3.00
4
ID R&D
91.86
10.36 (2)
97.48 (2)
2.99
5
K-IBS-DS
91.61
22.48 (6)
96.54 (5)
1.96
6
CTAI
90.59
21.88 (5)
96.40 (6)
-1.94
Sub-Task 2.3 (Overall Improvement): Synthetic + Real Data (Constrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
K-IBS-DS
95.42
9.49 (5)
98.14 (6)
-2.15
2
OPDAI
95.23
7.54 (1)
98.70 (1)
-0.52
3
CTAI
94.56
8.85 (4)
98.41 (3)
-6.01
4
CBSR-Samsung
94.20
8.62 (3)
98.17 (4)
-4.40
5
ADMIS
94.15
10.99 (6)
98.46 (2)
-1.10
6
ID R&D
94.05
8.00 (2)
98.16 (5)
0.07
Figure 12: DET curves of Sub-Tasks 1.1 (left), 1.2 (Middle), and 1.3 (right), including the results achieved for the top-1 team (i.e., ID R&D, ID R&D, ADMIS, respectively) in each demographic group.
V-BTask 2: Overall Improvement
Table VI provides the results achieved by participants in Task 2, focusing not only on bias mitigation but also other challenges in FR such as age, pose, and occlusions. Teams are ranked in descending order based on the average verification accuracy across the four databases. Notably, in all sub-tasks, the AVG is lower than the achieved in Task 1 for the BUPT-BalancedFace [43] database, showing the additional challenges introduced by the AgeDB [44], CFP-FP [45], and ROF [46] databases. This trend can also be observed in GAP results, which tend to be worse for Sub-Tasks 2.1 and 2.2 compared to Sub-Tasks 1.1 and 1.2, suggesting that it is far more difficult to emulate the conditions of these real databases with synthetic data by itself. For example, in Sub-Task 2.1, although the top-3 teams achieve high AVG results, they exhibit a considerable positive GAP value (i.e., OPDAI 91.93% AVG, 3.09% GAP; ID R&D 91.86% AVG, 2.99% GAP; ADMIS 91.19% AVG, 2.78% GAP), showing that a FR model trained only with real data (i.e., CASIA-WebFace) adapts better to adverse image conditions such as aging, pose, or occlusions. Focusing on Sub-Task 2.2, the team ranked top-1, i.e., Idiap-SynthDistill, achieves much better results compared to the best result of Sub-Task 2.1 (i.e., 93.50% vs. 91.93% AVG), proving that unlimited synthetic data can further improve the performance of the system. Finally, in Sub-Task 2.3, most teams report better AVG and higher negative GAP values (e.g., K-IBS-DS achieves 95.42% AVG, -2.15% GAP), proving again that synthetic data combined with real data can alleviate existing limitations within FR technology.
V-CDemographic Groups and Evaluation Databases
This section provides an in-depth analysis of the results in terms of the different demographic groups and individual databases considered in the 2 FRCSyn-onGoing. Figure 12 shows the Detection Error Tradeoff (DET) curves of Sub-Tasks 1.1, 1.2, and 1.3, including the results achieved for the top-1 team in each demographic group. For completeness, the information and graphical representations for all the teams can be found on the challenge Codalab platform121212https://codalab.lisn.upsaclay.fr/competitions/16970#results.
For Sub-Tasks 1.1 and 1.2, the team that achieves the first place, ID R&D, demonstrates high performance across the different demographic groups considered (above 96.50% Accuracy for all demographic groups). However, a slight gender bias can be observed (improvements of 1% between ‘Male’ and ‘Female’ labels for some ethnicities). Regarding the ethnicity, the proposed FR model showed better results for subjects from the ‘Indian’ ethnicity (99.20% Accuracy for Indian Male; 98.10% Accuracy for Indian Female). Finally, for Sub-Task 1.3, the winning team, ADMIS, also performs well across all demographic groups (all above 96% Accuracy). However, there exists variability in performance between different demographic groups. For example, ‘Asian Females’ and ‘Indian Females’ have the lowest Accuracy (96.20%) while ‘White Females’ have the highest Accuracy (98.60%).
Figure 13: DET curves of Sub-Tasks 2.1 (left), 2.2 (Middle), and 2.3 (right), including the results achieved for the top-1 team (i.e., OPDAI, Idiap-SynthDistill, K-IBS-DS, respectively) in each evaluation database.
Figure 13 shows the DET curves of Sub-Tasks 2.1, 2.2, and 2.3, including the results achieved for the top-1 team in each database. Analyzing the FR model proposed by the OPDAI team for Sub-Task 2.1, the spread of the curves indicates variability in the system performance across different databases, with the results from AgeDB (94.54% Accuracy) outperforming others. Moreover, in Sub-Task 2.2 the Idiap-SynthDistill FR model significantly improves the performance of Sub-Task 2.1 for AgeDB and CFP-FP databases (i.e., 96.72% and 96.14% Accuracy, respectively). Finally, for Sub-Task 2.3, the curves from the K-IBS-DS FR model are closely aligned for the AgeDB (96.89% Accuracy), BUPT (96.88% Accuracy), and CFP-FP (97.51% Accuracy), showing consistent and reliable performance across these databases. However, the curve from the ROF database remains the worst in each sub-task (i.e., 87.22%, 87.31%, and 90.39% Accuracy for Sub-Tasks 2.1, 2.2, and 2.3, respectively), reflecting that it is the most difficult database to emulate with synthetic data.
V-DPost-Challenge Analysis and Comparison with 1 Edition
Analyzing the contributions of all eleven top teams, we can observe the prevalence of well-established methodologies. Notably, most teams used DCFace [28] either independently or in conjunction with other synthetic databases such as GANDiffFace [21], DigiFace-1M [31], or IDiff-Face [29]. Furthermore, several teams, including CTAI and CBSR-Samsung, adopted interesting approaches involving synthetic data cleaning and selection. It is interesting to highlight the ID R&D and Idiap-SynthDistill teams as they considered novel methods to generate synthetic data. Specifically, the ID R&D team used an HDT [54] to generate synthetic facial images along with identity and style embeddings, wich were used by a StyleNAT [56] model to generate more variability in the synthetic data. Another example is the Idiap-SynthDistill team, which proposed an end-to-end method that dynamically generated facial images through StyleGAN2 [62] and trained a FR model through model distillation. Regarding the backbone architecture, all teams opted for either ResNet [36] or IResNet [48], mainly for their widespread adoption in state-of-the-art FR methodologies. The selection of the loss functions was also similar among the teams, with AdaFace [7] and ArcFace [6] the prevalent choices. Nevertheless, there were exceptions such as for the ID R&D team that used the recent UniFace [58], or the UNICA team that considered CosFace [61].
TABLE VII: Description of the best results achieved in the 1 and 2 FRCSyn-onGoing. Sub-Tasks 1.2 and 2.2 of the 2 FRCSyn-onGoing are not included in the table as they are novel sub-tasks only available in the 2 edition.
Sub-Task 1.1: Bias Mitigation
1 FRCSyn-onGoing
2 FRCSyn-onGoing
Team
TO [%]
GAP [%]
Team
TO [%]
GAP [%]
LENS
92.25
-0.74
ID R&D
96.73
-5.31
Sub-Task 1.3: Bias Mitigation
1 FRCSyn-onGoing
2 FRCSyn-onGoing
Team
TO [%]
GAP [%]
Team
TO [%]
GAP [%]
CBSR
95.25
-2.10
ADMIS
96.50
-1.33
Sub-Task 2.1: Overall Improvement
1 FRCSyn-onGoing
2 FRCSyn-onGoing
Team
AVG [%]
GAP [%]
Team
AVG [%]
GAP [%]
BOVIFOCR
90.50
2.66
OPDAI
91.93
3.09
Sub-Task 2.3: Overall Improvement
1 FRCSyn Challenge
2 FRCSyn-onGoing
Team
AVG [%]
GAP [%]
Team
AVG [%]
GAP [%]
BOVIFOCR
94.95
-3.69
K-IBS-DS
95.42
-2.15
Finally, we compare the results achieved in the 2 FRCSyn-onGoing with the results of the 1 edition [34]. Table VII shows the best results achieved in the 1 and 2 editions of the challenge, including also the GAP values. It is important to remark that Sub-Tasks 1.2 and 2.2 of the 2 FRCSyn-onGoing are not included in the analysis as they are novel sub-tasks only available in the 2 edition. Notably, two observations can be made: i) the main metric for ranking teams (i.e., TO and AVG) shows improvements across all cases in this 2 edition for both Task 1 (e.g., 96.73% vs. 92.25% TO in Sub-Task 1.1 and 96.50% vs. 95.25% TO in Sub-Task 1.3) and Task 2 (e.g., 91.93% vs. 90.50% AVG in Sub-Task 2.1 and 95.42% vs. 94.95% AVG in Sub-Task 2.3), and ii) in terms of the GAP value, the FR models of this 2 edition follow a similar trend compared to the 1st edition, achieving in most sub-tasks negative GAP values, remarking the benefits for training using synthetic data. In particular, for Sub-Task 1.1, a much higher negative GAP value is observed in this 2 edition (i.e., -5.31% vs. -0.74%). This result, together with the higher TO value, seems to be due to the generation of better synthetic data by the ID R&D team, with the proposal of novel generative methods, as indicated before. In addition, several conclusions can be drawn. First, the improvement in the main metric can be associated with the freedom to select the methodology to generate the synthetic data to train the FR models, as well as the application of data cleaning and selection techniques. Finally, the increasing GAP value can be associated also with the enhancement of the FR models, due to the proposal of different architectures and loss functions, as indicated before.
VI Conclusion
The 2 FRCSyn-onGoing has presented a comprehensive exploration of the applications of synthetic data in FR, effectively addressing existing limitations in the field. In this 2 edition, two additional sub-tasks have been introduced, showing that impressive results can be achieved using unlimited synthetic data, even outperforming in some cases the scenario of training with only real data. With an increased number of participants in this last edition, we have witnessed a considerable performance improvement in all sub-tasks in comparison to the 1 edition [34, 35]. This has been possible thanks to the proposal of novel methods to generate and select better synthetic data, as well as FR models and loss functions. These approaches can be compared across a variety of sub-tasks, with many being reproducible thanks to the materials made available by the participating teams.
Future studies will include recent AI techniques [80], to make sure that only the databases available by the challenge are used by participants. We will also perform a more detailed analysis of the results and comparison with recent challenges in the topic, such as SDFR [81], or evaluate over a more diverse set of databases that include other FR challenges, like quality, surveillance, or large distance images [82]. Finally, we plan to focus on the explainability of these FR models and the frameworks that generate synthetic images [8]. Debiasing face recognition models using synthetic datasets is an important task and in this challenge, we found that the use of synthetic data can furhter increase the performance. However, the concept of “bias” itself is complex and often subjective. What constitutes a “fair” representation can vary significantly depending on cultural context, individual experiences, and even personal beliefs. Therefore, debiasing efforts should be approached with an accurate understanding of the multifaceted nature of bias. Simply generating synthetic data to reflect a particular demographic distribution might not fully address the complexities of real-world inequalities. Rather than seeking to eliminate bias, perhaps a more productive approach is to pursue research in the direction of transparency, interpretability, and controllability. This translates to research questions that allow researchers to easily define what bias should be and allow them to fine-tune their models accordingly. Ultimately, the goal should be to develop face recognition systems that are not only accurate but also fair and ethical.
Acknowledgements
This study has received funding from the European Union’s Horizon 2020 TReSPAsS-ETN (No 860813) and is supported by INTER-ACTION (PID2021-126521OB-I00 MICINN/FEDER), Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), R&D Agreement DGGC/UAM/FUAM for Biometrics and Cybersecurity, and PowerAI+ (SI4/PJI/2024-00062, funded by Comunidad de Madrid through the grant agreement for the promotion of research and technology transfer at UAM).
It is also supported by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.
K-IBS-DS was supported by the Institute for Basic Science, Republic of Korea (IBS-R029-C2). UNICA-IGD-LSI was supported by the ARIS program P2-0250B.